You have two flashlights that have 0.500-A currents flowing through them. One flashlight has a single 1.50-V battery in its circuit, and the second flashlight has three 1.50-V batteries connected in a chain that provides 4.50 V. How much power is the battery in the first flashlight providing? How much power is each battery in the second flashlight providing?
Answer #1

power in the first flashlight

power P = I*V

I=0.5 A , V = 1.5 V

P = 0.5*1.5 = 0.75 W

power is each battery in the second flashlight

they are connected in seres

current constant I = 0.5 A and V = 1.5 V each one

power on each  flashlight = 0.5*1.5 = 0.75 W